1. A 10 percent increase in income causes the quantity of orange juice demanded to increase from 14,200 gallons to 18,800 gallons.  What is the income elasticity of demand for orange juice? (5 marks) 
  2. The cross elasticity between two goods is -2.8.  Provide examples of these two goods, and define the type of good.  Please ensure to provide an explanation to support your choice of type of good. (5 marks)
  3. Assume the country of Marvellous produces two services education (schools) and hospitals.  If it uses all of its resources – land, labour, capital and enterprise, it can produce a maximum of 40 hospitals or a maximum of 60 schools.  

a) Draw a production possibility curve illustrating the information provided above (7 marks)  


b) Using the production possibility sketched in part a explain how the following concepts are illustrated on the production possibility frontier

  1. Scarcity (1 mark)
  2. Choice (1 mark)
  3. Inefficiency (1 mark)
  4. Opportunity cost (2 marks)  

c) Illustrate the impact of the introduction of zoom technology, which would enhance the delivery of educational services, on the production possibility curve. (3 marks)


d) Illustrate the impact of an increase in the nation’s resources on the production possibility curve (3 marks) 




1
Expert's answer
2021-03-11T09:58:30-0500

(1) income elasticity

"\Delta Q/\Delta I"

"=(14200-18800)\/14200=-0.3"

"=0.3*100=30" %

"=-30\/10"

=-3

(2) complementary goods.

Bread and butter

When demand for bread Increase the demand for butter will increase proportionally.

Car and petrol.

When demand for cars Increase the demand for petrol will increase proportionally.

An increase in price of one item will leads to a decrease to the compliment because the demand of the main good decreases.

(3)(a)




(b)scarcity- The production possibility frontier illustrates scarcity by dividing production space into attainable and unattainable levels of production.

Choice-The production possibility frontier shows all possible combination of schools and hospitals available at a time.

Inefficiency-In the production possibility frontier, all points beneath the curve indicates inefficient.

Opportunity cost-The production possibility frontier slope indicates the opportunity cost of producing schools versus hospitals.

(c) Introduction of zoom technology will shift the production possibility curve to the right or outward.

(d) Increase of national resources will shift the production possibility curve to the right or outward.